Have We Seen the Top Quark?

Pekka K. Sinervo University of Toronto / IPP (representing the CDF Collaboration)

1. Why Do We Need to Find Top?

2. How Do We Look for It?

3. The CDF Experiment

4. Search for Dilepton Events

5. Search for Lepton+Jet Events

6. How Heavy is Top?

7. Conclusions

University of Waterloo May 10, 1994